How many 9-bit strings (that is, bit strings of length 9) are there which:



(a) Start with the sub-string 101? Explain.



(b) Have weight 5 (i.e., contain exactly five 1’s) and start with the sub-string 101? Explain.



(c) Either start with 101 or end with 11 (or both)? Explain.



(d) Have weight 5 and either start with 101 or end with 11 (or both)?Explain

1
Expert's answer
2022-03-08T08:34:01-0500

Solution (a)

There are 9-bits, and we start with 101. It means there are 6 remaining bits.

Therefore, there will be "2^6 = 64" strings (9-bits strings) 


Solution (b)

Given the weight of the string is 5 and since it starts with 101, and has a total of 5 1’s, it means the last 6-bits must have three 1’s.

So, there will be 6 places out of which three will be occupied by three 1’s. So, there will be "6C3", strings (9-bits string) starting with 101 and will have weight equal to 5. 




Solution (c)

Let us consider the string starts with 101.

Then there will be "2^6 = 64" strings (9-bits strings)

 

Let us consider the string ends with 11.

Then there will be "2^7 = 128" strings (9-bits strings)

 

When the string starts with 101 and ends with 11, then there will be "2^4 = 16" strings (9-bits strings) 


Solution (d)

 

If the 9-bits string has a weight of 5, then

From the 9 places to have digits, there are 5 places to have 1's.

Therefore, there will be 9C5 = 126 strings (9-bits strings)

n(A) = 9C5 = 126 strings (9-bits strings)

  

The string starts with 101 and has a weight of 5. Then there will be 6C3 =20 strings (9-bits strings)

That is 



Then there will be 20 strings (9-bits strings)

 

Now when the strings start with 101, ends with 11 and have the weight 5, we will have

"n(A or B) = n(A) + n(B) – n(A and B)\\"


"n(A or B) = 126 + 64 \u2013 20"


"n(A or B) = 170"